# Applications of Spiking Neural Networks in Visual Place Recognition

### v3

@@ IV-E Baseline Methods
Non-modular SNN [59]: is a single three-layer SNN model that is enlarged to accommodate for learning a significantly larger number of places. The model is trained on the reference dataset, and tested on the query dataset. This comparison is included to show the scalability advantages of ensembling and modularity in terms of both performance and computational time as the number of places to learn increases.
@@ V-C Comparison of Ensemble of Modular SNNs with a Sequence Matcher to Conventional VPR Techniques
V-D An Indicator for Sequence Matching Responsiveness
@@ V-E Ensembling: How Much Does It Help?
This section evaluates the effect of ensembling on our Modular SNN, and provides comparisons to GCL and AP-GeM ensembles. The ensemble members in the case of our Modular SNN are homogeneous as they share the same network architecture and training data; their differences lie in the random initialization values of weights and the random sequence of input images, as elaborated in Section III-D. The GCL ensembles are also homogeneous with consistent network architecture and training datasets and only differing in random initialization of weights and order of input images. The AP-GeM ensembles are heterogeneous, showing diverse Convolutional Neural Network (CNN) backbone and/or training datasets. The three architectures are a ResNet50 and a ResNet101, both trained on the Landmarks-clean dataset, and a Res101 trained on the Google-Landmarks Dataset, as described in Section IV-E.
@@ V-F Ablation Study on Member Randomization in Ensemble of Modular SNNs
This section provides an ablation study on the R@1 performance of our Ensemble of Modular SNNs with and without input image order shuffling, and with and without different random weight initialization that is applied to the ensemble members. In our previous conference paper [61], we provided consecutive input images of a traverse to the modules for training. We initialized the weights of all modules using the same random values. Here, to increase the diversity among the ensemble members, we instead shuffled the reference images of each traverse for the training process, and then fed these shuffled images to the modules (Figure 1). Moreover, we initialized the weights of each member, Modular SNN, using different random values, while within each member, using the same set of random values for all modules.
@@
Table II presents the R@1 performance of these four ensemble variants on the Oxford RobotCar dataset, where each ensemble contains three ensemble members. The first variant is an ensemble model where no randomization is applied, resulting in identical ensemble members and thus mirroring the performance of a single member, which is 10.4%. The ensemble members in the second variant differ only in their initial random weights, resulting in a R@1 of 11.0%. The third variant varies the ensemble members with only in the shuffled order of images, yielding in a R@1 performance of 19.1%. Lastly, the fourth variant combines both randomization of weights and shuffled order of images, producing the highest R@1 performance of 19.8%.
Excluding the first variant, which negates the ensembling effect, the remaining three variants show similar R@1 performance improvements over their mean R@1 member performances. Randomizing the shuffled order of images significantly enhances ensemble performance compared to just randomizing the weights. The combination of both strategies obtains the highest R@1, albeit with a marginal improvement over the third variant, where only shuffled order of images are randomized.
@@ VI Discussions and Conclusions
Our research illuminates the significant potential of SNNs for robotic navigation, presenting a solution that is scalable, and robust for place recognition tasks. Our SNN-based approach is particularly responsive to ensembling and sequence matching techniques, as evidenced by its performance increase when these techniques are applied. These techniques significantly enhance its robustness to high appearance changes, and its generalization ability across diverse environments. These advancements in SNN technology not only enhance the efficiency of robotic navigation systems but also have vast applicability across various real-world robotics applications. Our findings are particularly promising for resource-constrained robots, such as those deployed in challenging environments including space and underwater, where the focus on edge computing and considerations for SWaP (Size, Weight, and Power) emphasize its suitability for these rigorous settings.

### v4

@@ IV-E8 Non-modular SNN [59]
is a single three-layer SNN model that is enlarged to accommodate for learning a significantly larger number of places. The model is trained on the reference dataset, and tested on the query dataset. This comparison is included to show the scalability advantages of ensembling and modularity in terms of both performance and computational time as the number of places to learn increases.
@@ V-C Comparison of Ensemble of Modular SNNs with a Sequence Matcher to Conventional VPR Techniques
V-D Indicator for Sequence Matching Responsiveness
@@ V-E Ensembling: How Much Does It Help?
This section evaluates the effect of ensembling on our Modular SNN, and provides comparisons to GCL and AP-GeM ensembles. The ensemble members in the case of our Modular SNN are homogeneous as they share the same network architecture and training data; their differences lie in the random initialization values of weights and the random sequence of input images, as elaborated in Section III-D. The GCL ensembles are also homogeneous with consistent network architecture and training datasets and only differing in random initialization of weights and order of input images. The AP-GeM ensembles are heterogeneous, showing diverse Convolutional Neural Network (CNN) backbone and/or training datasets. The three architectures are a ResNet50 and a ResNet101, both trained on the Landmarks-clean dataset, and a ResNet101 trained on the Google-Landmarks Dataset, as described in Section IV-E.
@@ V-F Ablation Study on Member Randomization in Ensemble of Modular SNNs
This section provides an ablation study on the R@1 performance of our Ensemble of Modular SNNs with and without input image order shuffling, and with and without different random weight initialization that is applied to the ensemble members. In our previous conference paper [61], we provided consecutive input images of a traverse to the modules for training. We initialized the weights of all modules using the same random values. Here, to increase the diversity among the ensemble members, we instead shuffled the reference images of each traverse for the training process, and then fed these shuffled images to the modules (Figure 1). Moreover, we initialized the weights of each member, Modular SNN, using different random values, while within each member, using the same set of random values for all modules. Table II presents the R@1 performance of these four ensemble variants on the Oxford RobotCar dataset, where each ensemble contains three ensemble members. The first variant is an ensemble model where no randomization is applied, resulting in identical ensemble members and thus mirroring the performance of a single member, which is 10.4%. The ensemble members in the second variant differ only in their initial random weights, resulting in a R@1 of 11.0%. The third variant varies the ensemble members with only in the shuffled order of images, yielding in a R@1 performance of 19.1%. Lastly, the fourth variant combines both randomization of weights and shuffled order of images, producing the highest R@1 performance of 19.8%. Excluding the first variant, which negates the ensembling effect, the remaining three variants show similar R@1 performance improvements over their mean R@1 member performances. Randomizing the shuffled order of images significantly enhances ensemble performance compared to just randomizing the weights. The combination of both strategies obtains the highest R@1, albeit with a marginal improvement over the third variant, where only shuffled order of images are randomized.
@@ VI Conclusion
Our research illuminates the significant potential of SNNs for robotic navigation, presenting a solution that is scalable, and robust for place recognition tasks. Our SNN-based approach is particularly responsive to ensembling and sequence matching techniques, as evidenced by its performance increase when these techniques are applied. These techniques significantly enhance its robustness to high appearance changes, and its generalization ability across diverse environments. These advancements in SNN technology not only enhance the efficiency of robotic navigation systems but also have vast applicability across various real-world robotics applications. Our findings are particularly promising for resource-constrained robots, such as those deployed in challenging environments including space and underwater, where the focus on edge computing and considerations for size, weight, and power emphasize its suitability for these rigorous settings.
VII Acknowledgment
The authors would like to thank the Queensland University of Technology (QUT) for continued support through the Centre for Robotics. The authors would also like to thank Dr. A. Hines, T. Joseph, Dr. C. Malone, and G. B. Nair for their valuable insights on the drafts of this manuscript, and the QUT eResearch services for providing computational resources via the QUT High Performance Computing system.